# Warehouse storage and retrieval optimization via clustering, dynamic systems modeling, and GPU-accelerated routing

### v1

@@ 5. Discussion and conclusions
When we performed experiments of using unsupervised clustering to reduce computational complexity when determining routes as described at the end of Section 4, we often found that the computed route coincided with the optimal route after 300 iterations (as well as at earlier stages). However, since this is not guaranteed, we deliberately chose to present an experiment where this was not the case – the middle and right panels of Figure 9 are indeed different. In addition, when the separation between the different regions of stopping nodes (cf. the dashed rectangles in Figure 2) varies, some scenarios might call for a natural collapsing of two (perhaps somewhat overlapping) regions into one, thereby changing the clustering scheme. There is potential for analyzing the effects of such phenomena as well, but that is beyond the scope of this paper.
@@
The implementation of the Bellman-Ford algorithm in CUDA C demonstrates that GPU-based parallelism is highly effective for solving routing problems, particularly when evaluating large numbers of path combinations. While the exponential growth of permutations presents a challenge, especially as the number of nodes increases, this limitation can be mitigated by dividing large routing problems into smaller sub-routes. By processing these segments independently, it is possible to maintain high computational efficiency without exceeding the GPU’s memory or thread capacity. This segmentation strategy ensures that GPU acceleration remains a scalable and practical solution even for complex graph traversal problems. Therefore, the combination of Bellman-Ford and CUDA C provides a flexible framework that balances algorithmic precision with hardware-aware execution.
The application of the mathematical models and numerical methods presented in this paper closely parallels contemporary advancements in industry practice. In particular, the dynamic clustering techniques described in our numerical experiments, where warehouse storage layouts adapt continuously based on changing order patterns, are exemplified in operations such as those in Amazon’s fulfillment centers. Amazon employs sophisticated robotic picking algorithms, dynamically optimizing the selection and clustering of inventory units to enhance operational efficiency and reduce picking times [1]. This real-world application aligns closely with our mathematical formulation of order clustering and its resulting impact on route optimization, underscoring the practical viability of our proposed approach. Our numerical experiments show that while random variability in orders reduces the clarity of clustering effects on the definition and refinement of picking regions, the overall impact remains significant. This suggests that our approach is robust enough for practical deployment in real-world environments characterized by inherent order variability.
Our exploration of GPU-accelerated route optimization algorithms, in particular the Bellman-Ford algorithm implemented through CUDA-based parallelization, also finds a clear reflection in industry. Alibaba’s implementation of vehicle routing algorithms, optimized through GPU computing techniques, exemplifies the practical benefits of this computational strategy, which has significantly enhanced logistics performance and reduced operational costs through improved route optimization [12]. The success of these implementations in a commercial context validates the computational approach proposed in our numerical experiments, demonstrating the scalability and effectiveness of GPU-based solutions for complex warehouse routing scenarios.
@@ Appendix A WMS loop
The Order Processing algorithm converts an order ID into a string representation and determines the position to test. It then updates the stock based on the order string and increments the pick count in the graph $G$.
@@
Function ProcessOrder($order\_id,ordergraf$):
@@
       UpdateStock($G,order\_string$)
       $increment\_pick\_count(G,order\_string)$
@@
The Update Stock algorithm updates the stock in the graph $G$ based on the provided order string. It ensures that the inventory levels are adjusted according to the orders processed.
@@
Function UpdateStock($G,order\_string$):
       $update\_stock\_from\_order(G,order\_string)$
@@
The Check Stock algorithm verifies if there is sufficient stock in the graph $G$ to fulfill the next order string. It returns a Boolean value indicating whether the stock is sufficient.
@@
Function CheckStock($G,next\_order\_string$):
       return $is\_sufficient\_stock\_for\_order(G,next\_order\_string)$
@@
The Check Stock and Move Articles algorithm first checks if there is sufficient stock for the next order. If the stock is insufficient, it identifies empty nodes and calculates the distances to determine the optimal movement of articles to balance the stock.
@@
Function CheckStockAndMove($G,ordergraf,order\_id$):
@@
      if CheckStock($G,next\_order\_string$) = False then
@@
            for $i\leftarrow 0$ to $length(distance2)-1$ do
@@
                  if $test\_pos$ then
                         MoveArticle($G,distance2[i][0],min\_distance[0],blocked\_orders[i][3]$)
@@
The Main Order Processing Algorithm coordinates the overall process of handling orders. It iterates through a specified number of repetitions and processes each order by converting it to a string, updating the stock, and checking the stock to move articles as necessary.
@@
       for $repetition\leftarrow 0$ to $n\_repetitioner-1$ do
             for $order\_id\leftarrow 1$ to $order\_index$ do
                   ProcessOrder($order\_id,ordergraf$)
                  if $order\_id<number\_of\_orders$ then
                         CheckStockAndMove($G,ordergraf,order\_id,order\_string$)

### v2

@@ 6. Discussion and conclusions
The results presented above demonstrate that the proposed clustering–RDS framework can reorganize warehouse layouts dynamically, yielding substantial improvements in route efficiency and cluster compactness. While the experiments rely on synthetic but representative warehouse configurations, the observed effects are systematic and robust across repetitions. Although the simulation setting abstracts from several operational complexities, it captures the essential structural dynamics of storage-location assignment and product reorganization.
The following discussion outlines the robustness and limitations of the proposed framework, evaluates its practical relevance, and identifies directions for future research and industrial deployment.
@@
The implementation of the Bellman–Ford algorithm in CUDA C demonstrates that GPU-based parallelism is highly effective for solving routing problems, particularly when evaluating large numbers of path combinations. While the exponential growth of permutations presents a challenge, especially as the number of nodes increases, this limitation can be mitigated by dividing large routing problems into smaller sub-routes. By processing these segments independently, it is possible to maintain high computational efficiency without exceeding the GPU’s memory or thread capacity. This segmentation strategy ensures that GPU acceleration remains a scalable and practical solution even for complex graph traversal problems. Therefore, the combination of Bellman–Ford and CUDA C provides a flexible framework that balances algorithmic precision with hardware-aware execution.
The numerical experiments demonstrate that the proposed clustering–RDS framework remains robust under a wide range of order variability and warehouse dynamics. In smaller-scale experiments, performance declines as stochastic noise increases, yet clustering effects remain statistically significant and clearly measurable through both geometric and silhouette-based indicators. In the large-scale setting with 100 000100\,000 nodes, the method exhibits strong improvements in cluster compactness and separation across all noise levels, showing that its core mechanisms scale effectively even when orders occupy only a small fraction of total capacity. This robustness arises from the adaptive interaction between clustering and relocation: as orders fluctuate, the RDS formulation continuously rebalances the spatial distribution of products to maintain compact storage regions. Such behavior mirrors the adaptive, self-organizing tendencies of modern fulfilment systems, in which storage assignments evolve dynamically in response to changing demand.
While the framework captures key structural dynamics of warehouse reorganization, it relies on several simplifying assumptions that delimit its immediate operational realism. In particular, the model currently assumes single-SKU storage locations (one article type per node), and that articles may be relocated freely at each iteration. These abstractions allow the effects of clustering and route optimization to be studied in isolation, providing a clear baseline for understanding how spatial reorganization improves efficiency. In practice, relocations are constrained by available labour and may involve non-negligible delays. Such constraints can be incorporated in at least two ways: by placing explicit limits on the number of relocations per iteration (labour capacity constraints), or by assigning a relocation cost or delay to each move (time-based constraints). As discussed in Section 2.1, the mathematical formulation can be extended to incorporate multi-SKU nodes, which together with extensions such as relocation costs and time-based capacity limits represent a natural next step toward operational deployment of the model in real fulfilment environments.
In addition, replenishment—the restocking of depleted picking locations from reserve storage—is implicitly represented in our model through the relocation mechanism itself: when inventory of a given article type at a node is exhausted, the model replenishes it dynamically by assigning a new storage location based on the current clustering configuration. This can be interpreted as a form of adaptive, demand-driven replenishment rather than a static restocking process. Alternatively, replenishment operations could be modeled explicitly as delayed or capacity-limited relocation events, which would constitute natural extensions of the present model and could be integrated with our clustering framework in future work.
Another direction of future research concerns energy-aware warehouse operations, where sustainable storage assignment and system design are increasingly studied in the context of automated systems [28]; such considerations could in principle be combined with clustering-based reorganization to jointly optimize efficiency and sustainability.
Comparable principles are already applied in industry. For instance, Amazon’s robotic picking algorithms dynamically re-cluster storage units based on changing order patterns [1], while Alibaba’s GPU-accelerated routing systems achieve real-time optimization at large scale [22]. These examples demonstrate that the computational and organizational mechanisms analyzed here align with successful industrial practices, supporting the practical viability of the proposed framework.
Future research should also extend the framework toward operational deployment by embedding labour and equipment constraints directly into the RDS model through capacity or cost terms, enabling realistic scheduling of human or robotic pickers. Further, coupling the model with digital-twin or WMS simulation platforms would facilitate real-time decision support, enabling policy evaluation and benchmarking. Empirical validation on real warehouse datasets—beyond the synthetic but representative instances studied here—would establish quantitative benchmarks for cost savings and throughput gains, paving the way for practical adoption in large-scale industrial settings.
Managerial implications and insights
From a practical standpoint, these findings translate into several actionable insights for warehouse managers. The proposed framework serves as a strategic decision-support tool for warehouse planners, logistics engineers, and managers responsible for optimizing storage layouts and order-picking efficiency. By integrating dynamic clustering with route optimization, it provides quantitative insight into how reorganizing stock locations can improve picking performance as demand patterns evolve. The model can be embedded into a WMS platform or used in an offline planning environment to evaluate alternative zoning and replenishment strategies before physical implementation, thereby identifying configurations that minimize travel distances and enhance space utilization.
The numerical experiments demonstrate that the clustering-based reorganization substantially improves order consolidation and reduces average route length, with the approximate GPU-accelerated routing procedure yielding paths within 2–14% of the optimal solution while maintaining sub-second runtimes for problem sizes up to $10^{5}$ nodes. These results indicate that dynamic, data-driven reallocation of stock locations—implemented even periodically rather than continuously—can produce meaningful operational gains without major infrastructure changes. Compared to traditional static ABC zoning, the proposed method adapts automatically to shifting order profiles, maintaining compact and well-separated product clusters even under moderate noise conditions.
From a managerial perspective, the framework’s simplicity and modularity make it well suited for exploratory “what-if” analyses in digital-twin environments or as an analytical layer within existing WMS platforms. Implementation requires only basic positional data and historical order information, making it both accessible and interpretable for practitioners. In practice, further calibration would be needed to incorporate labour availability, relocation costs, and replenishment schedules—elements that can be introduced as constraints or penalty terms in future extensions. Overall, the framework provides a scalable, computationally efficient foundation for developing adaptive and energy-aware storage assignment policies in next-generation smart warehouses. Taken together, these results demonstrate how mathematically grounded models can directly inform data-driven warehouse management, bridging analytical rigor and managerial practice.
@@ Appendix A WMS loop
For clarity, Table 5 summarizes the notation used in Algorithms 15.
Table 5. Notation used in Algorithms 15.
Symbol / Function Description
$G$ Warehouse graph containing product positions and balances
order_idorder\_id Identifier of the current order being processed
ordergrafordergraf Order graph containing article numbers and ordered quantities
order_indexorder\_index Total number of orders in the current batch
n_repetitionern\_repetitioner Number of repetitions (iterations) of the main loop
order_stringorder\_string Encoded string representation of an order
test_postest\_pos Position index used for testing or stock updates
next_order_stringnext\_order\_string Encoded representation of the subsequent order
empty_nodesempty\_nodes Set of empty warehouse nodes available for relocation
blocked_ordersblocked\_orders Set of orders blocked due to insufficient stock
specific_pointspecific\_point Cluster center associated with the current order
distancedistance Distance list between empty nodes and cluster center
$distance2$ Distance list between blocked orders and cluster center
min_distancemin\_distance Minimum distance among candidate relocation nodes
ProcessOrder() Converts an order ID into a string and updates stock levels
UpdateStock() Updates inventory in $G$ based on the processed order
CheckStock() Returns True if sufficient stock exists for an order
FindEmptyNodes() Identifies empty nodes available for article relocation
CalculateDistance() Computes distances between node sets and a reference point
MoveArticle() Moves an article from one node to another
CheckStockAndMove() Combines stock checking and relocation to balance inventory
Main() Main WMS loop coordinating order processing and relocation
The Order Processing algorithm converts an order ID into a string representation and determines the position to test. It then updates the stock based on the order string and increments the pick count in the graph $G$, see Algorithm 1.
@@
Function ProcessOrder(order_id,ordergraforder\_id,ordergraf):
@@
 UpdateStock(G,order_stringG,order\_string)
 increment_pick_count(G,order_string)increment\_pick\_count(G,order\_string)
@@
The Update Stock algorithm updates the stock in the graph $G$ based on the provided order string. It ensures that the inventory levels are adjusted according to the orders processed, see Algorithm 2.
@@
Function UpdateStock(G,order_stringG,order\_string):
 update_stock_from_order(G,order_string)update\_stock\_from\_order(G,order\_string)
@@
The Check Stock algorithm verifies if there is sufficient stock in the graph $G$ to fulfill the next order string. It returns a Boolean value indicating whether the stock is sufficient, see Algorithm 3.
@@
Function CheckStock(G,next_order_stringG,next\_order\_string):
 return is_sufficient_stock_for_order(G,next_order_string)is\_sufficient\_stock\_for\_order(G,next\_order\_string)
@@
The Check Stock and Move Articles algorithm first checks if there is sufficient stock for the next order. If the stock is insufficient, it identifies empty nodes and calculates the distances to determine the optimal movement of articles to balance the stock, see Algorithm 4.
@@
Function CheckStockAndMove(G,ordergraf,order_idG,ordergraf,order\_id):
@@
 if CheckStock(G,next_order_stringG,next\_order\_string) = False then
@@
    for $i\leftarrow 0$ to length(distance2)1length(distance2)-1 do
@@
       if test_postest\_pos then
          MoveArticle(G,distance2[i][0],min_distance[0],blocked_orders[i][3]G,distance2[i][0],min\_distance[0],blocked\_orders[i][3])
@@
The Main Order Processing Algorithm coordinates the overall process of handling orders. It iterates through a specified number of repetitions and processes each order by converting it to a string, updating the stock, and checking the stock to move articles as necessary, see Algorithm 5.
@@
 for $repetition\leftarrow 0$ to n_repetitioner1n\_repetitioner-1 do
    for $order\_id\leftarrow 1$ to order_indexorder\_index do
       ProcessOrder(order_id,ordergraforder\_id,ordergraf)
       if order_id<number_of_ordersorder\_id<number\_of\_orders then
          CheckStockAndMove(G,ordergraf,order_id,order_stringG,ordergraf,order\_id,order\_string)